OpenStudy (anonymous):

Which arrangement for these numbers is from least to greatest? 3/5, 5/8, 6/7

OpenStudy (anonymous):

3/5 , 5/8 , 6/7

OpenStudy (tkhunny):

Give the decimal equivalents a quick look.

OpenStudy (anonymous):

I would assume 3/5, 5/8, 6/7 are right because.. 3/5 = 0.6 5/8 = 0.625 6/7 = 0.85714286 Or am I wrong?

OpenStudy (tkhunny):

Why do you doubt? It's hard to go wrong when you PROVE IT - like you just did.